# GravSphere2: A higher-order Jeans method for mass-modeling spherical stellar systems

### v1

@@ 1 Introduction
Many stellar systems are approximately spherical, including globular clusters (GCs) and nearby dwarf spheroidal galaxies (dSphs). Both GCs and dSphs contain seemingly invisible mass components that can be inferred from the anomalous motions of their stars and gas (see e.g. Bosma 1978; Rubin et al. 1980; Trimble 1987; Łokas 2002; Van der Marel & Anderson 2010; Battaglia et al. 2013, for more recent examples in GCs and dSphs (respectively), see e.g. Evans et al. 2022; Dickson et al. 2024; Smith et al. 2024a; Bañares-Hernández et al. 2025 and Read et al. 2019; Hayashi et al. 2020; Battaglia & Nipoti 2022; Bañares-Hernández et al. 2023; De Leo et al. 2024; Arroyo-Polonio et al. 2025; Pascale et al. 2025; Vitral et al. 2025). For GCs, the invisible mass most likely constitutes stellar remnants or an ‘intermediate-mass black hole’ (IMBH), though some GCs may also contain dark matter. By contrast, dSphs appear to be dominated at all radii by dark matter. With increasingly available spectroscopic and astrometric data from recent surveys, GCs promise an unprecedented natural laboratory to study stellar remnants and hunt for IMBHs (Mezcua, 2017; Neumayer et al., 2020; Greene et al., 2020; Askar et al., 2023; Vitral et al., 2023; Häberle et al., 2024; Bañares-Hernández et al., 2025), while dSphs promise the same for dark matter (Boehm et al., 2017; Read et al., 2018, 2019; Battaglia & Nipoti, 2022; De Leo et al., 2024; Vitral et al., 2024, 2025; Zimmermann et al., 2025).
The above proliferation in data has led to an associated proliferation in mass-modeling methodologies. All methods seek to solve the steady-state Collisionless Boltzmann Equation (CBE), also known as Vlasov Equation (e.g. Binney & Tremaine 2008):
@@
By solving for the gravitational potential and subtracting the expected potential from all visible components we can infer, and study, the presence of invisible masses like black holes and dark matter (e.g. Van der Marel & Anderson 2010; Walker et al. 2009), which is usually done in the steady state approximation, assuming $\frac{\partial f}{\partial t}=0$.
@@
There are number of methods for solving the CBE. Jeans methods solve moments of the CBE (Jeans, 1922), having the advantages that they are fast [^1] This is particularly timely when assessing increasingly large samples with thousands (in some cases approaching millions, Häberle et al. 2025) of tracers. and minimize potential biases that could creep in due to assumptions about the form of the distribution function, $f$ (e.g. Cappellari 2008; Watkins et al. 2013; Mamon et al. 2013; Cappellari 2015; Read & Steger 2017). The main disadvantages, however, are that the data typically need to be binned, while $f$ is not determined. This leads to some loss of information and makes it hard to incorporate stars of uncertain membership, binaries and/or rigorously model selection effects (e.g. Read et al., 2021). Furthermore, since $f$ is not determined, a physical $f$ is not guaranteed. This can result in models that formally require a negative $f$ in some regions of phase space (Dejonghe & Merritt, 1992; Richardson & Fairbairn, 2013) and that should, therefore, be discarded (e.g. An & Evans, 2006).
@@
Distribution function methods assume some form for $f$ and then fit this to the data (e.g. Gieles & Zocchi 2015; Vasiliev 2019; Read et al. 2021; Dickson et al. 2023). $f$ can also be implicitly expressed as a superposition of orbits with ‘Schwarzschild methods’ (e.g. Schwarzschild 1979; Chanamé et al. 2008; van den Bosch et al. 2008; Breddels et al. 2013) or with ‘made-to-measure’ methods, in which $N$-body simulations are evolved to fit some data constraints (e.g. Syer & Tremaine 1996; de Lorenzi et al. 2007; Dehnen 2009; Hunt & Kawata 2013). We also note non-parametric distribution function methods, such as the one from Li et al. (2025). Distribution function methods have the advantages that no data binning is required and $f$ is determined by the data. However, if the functional form for $f$ (or its implicit specification via other methods) does not encompass the true solution, then bias can creep in.
@@
Finally, numerous machine-learning methods have been proposed that learn from training data to either accelerate distribution function fitting or incorporate ‘cosmological priors’ into the mass-modeling pipeline (e.g. Expósito-Márquez et al. 2023; Nguyen et al. 2025; Lim et al. 2025; Sarrato-Alós et al. 2025).
@@
Given the proliferation of mass-modeling methods, we may wonder why yet another method is needed. However, there are two key problems that remain to be addressed, even for spherical stellar systems. Firstly, while the amount and quality of data continue to improve, there will always be interesting systems with very few stellar tracers (e.g. Collins et al. 2021; Smith et al. 2024b; Pickett et al. 2025). Typically, simple mass estimators are used for low tracer numbers, but these have been shown to be potentially quite biased (Laporte et al., 2019). More sophisticated, unbiased, mass-modeling methods for low tracer numbers are only just being developed (e.g. Nguyen et al. 2025).
@@
Secondly, there is a well-known degeneracy between the cumulative mass (which we are typically most interested in) and the velocity anisotropy profile – a measure of how circular or radial the orbit distribution is (e.g. Dejonghe 1987; Merritt 1987; Merrifield & Kent 1990; Wilkinson et al. 2002; Łokas & Mamon 2003; de Lorenzi et al. 2009; Read & Steger 2017; Read et al. 2021). For Jeans modeling, this occurs because the Jeans equations do not constitute a closed system, meaning that for a given mass model and tracer profile, a unique solution is not guaranteed (e.g. Binney & Tremaine 2008). This typically arises when only projected line-of-sight (LOS) velocities are available.
@@
Two key methods have been proposed to break the mass-anisotropy degeneracy. Firstly, one can incorporate higher moments of the velocity distribution as data constraints. At infinite order, these fully specify the phase-space distribution function and ensure its positivity (Merrifield & Kent, 1990; Dejonghe & Merritt, 1992; Richardson & Fairbairn, 2013). After the usually-used second moment, the next non-trivial moment for a spherical, non-rotating, pressure-supported system is the fourth moment. This encodes information about deviations from Gaussianity, including the ‘tailedness’ of the distribution and the kurtosis. It can be incorporated directly into the Jeans equations, thereby introducing a new fourth-order anisotropy parameter (Merrifield & Kent, 1990; Łokas & Mamon, 2003; Łokas et al., 2005; Richardson & Fairbairn, 2013; Wardana et al., 2025). Alternatively, one can perform luminosity-weighted integrals over the fourth moment to derive ‘Virial Shape Parameters’ (VSPs) that marginalize out the higher-order anisotropy (Merrifield & Kent, 1990; Richardson & Fairbairn, 2014; Read & Steger, 2017). Secondly, one can include proper motions (PMs) to gain additional phase-space information that is independent of the line-of-sight velocities (e.g. Van der Marel & Anderson 2010; Watkins et al. 2013). Until recently, such PMs have only been available for GCs (e.g. McNamara et al. 2003; Anderson & Van der Marel 2010; Vitral et al. 2022; Dickson et al. 2023; Häberle et al. 2024; Bañares-Hernández et al. 2025). However, data are becoming increasingly available also for nearby dwarf galaxies, with recent mass models incorporating PMs having been recently reported for the Draco and Sculptor dSphs (Vitral et al., 2024, 2025).
@@
Although Jeans mass-modeling methods typically use data binning (e.g. Read & Steger, 2017), this is not actually a requirement. Several authors have explored assuming a local Gaussian distribution function constrained by the second moment to create a bin-free Jeans method (e.g. Watkins et al. 2013; Mamon et al. 2013). With the widespread availability of individually resolved stellar velocities, this approach is preferable and has all the advantages of having $f$ that distribution function methods enjoy. [^2] This is true, up to a given order of moments, provided that the Jeans method has those moments self-consistently implemented in the velocity distribution function.
@@
In this work, we present the GravSphere2 mass-modeling method. This builds on GravSphere(Read & Steger, 2017; Collins et al., 2021) with the specific goals of: (i) reducing bias; (ii) maximizing the information content in the data; (iii) avoiding the need for data binning; and (iv) being able to successfully model, without bias, very low numbers of tracers. To achieve these goals, we self-consistently solve the second and fourth-order Jeans equations, following the framework in Richardson & Fairbairn (2013) and extending it to include a new observable: fourth-order PMs. We show that this breaks degeneracies in the higher-order Jeans equations by solving along their three projected components. Unlike previous approaches, which have made more restrictive assumptions (e.g. Merrifield & Kent, 1990; Łokas & Mamon, 2003; Łokas et al., 2005; Richardson & Fairbairn, 2013; Wardana et al., 2025), we also include fully independent and radially varying anisotropies at both orders. We model individual stellar velocities, rather than binning the data, by introducing a local non-Gaussian velocity probability distribution function (PDF) constrained by the second and fourth moments. We also model the positions of stars without binning, where data are available. We test these improvements using mock data for dSphs and we make the Python package, GravSphere2, publicly available at: https://github.com/dadams42/GravSphere2.
@@
This paper is organized as follows. In Sec. 2, we introduce the modeling methodology behind GravSphere2, including our method for solving the fourth-order Jeans equations, which we apply following Richardson & Fairbairn (2013) and extend to include PMs. We describe our implementation using non-Gaussian local velocity PDFs for individual tracers, following Sanders & Evans 2020, in Appendix A, and the functional forms we assume for our mass model and the second and fourth-order stellar velocity anisotropy profiles in Appendix B. In Sec. 3, we introduce the mock data which we use to test GravSphere2. In Appendix C, we describe how we integrate the approach introduced in Sec. 2 to fit these data, specifying the form of the likelihood function and our assumed priors. Sec. 4 shows the results of our tests of GravSphere2 on the mock data. We compare these results with previous methods that have been applied to these same data in the literature, including previous versions of GravSphere. We also explore the limit of low tracer numbers (i.e. corresponding to the faintest or most distant systems) and we compare the performance of GravSphere2 with simple mass estimators that are widely used in the literature. In Sec. 5, we briefly discuss the significance of this work and we add some concluding remarks, highlighting some upcoming works that use the new GravSphere2 method.
@@ 2 Modeling stellar kinematics with GravSphere2
GravSphere is a numerical Jeans solver that was originally introduced in Read & Steger (2017). It has since been further extended and extensively tested with mock data (Read & Steger, 2017; Read et al., 2018; Gregory et al., 2019; Genina et al., 2020; Read et al., 2021; Collins et al., 2021; Genina et al., 2022; De Leo et al., 2024; Nguyen et al., 2025; Tchiorniy & Genina, 2025) including simulations deliberately chosen to explore expected bias in dynamical models due to breaking of key assumptions, such as spherical symmetry (triaxial mocks) and dynamical pseudo-equilibrium (tidally stripped mocks). A public version of this GravSphere1 code, called pyGravSphere, was introduced in Genina et al. (2020). [^3] Available at: https://github.com/AnnaGenina/pyGravSphere. Its most recent version was introduced and made publicly available in Collins et al. (2021) [^4] Available at: https://github.com/justinread/gravsphere. which improved the velocity binning with the binulator feature, improved the mass model and improved the way that the Virial Shape Parameters contributed to the likelihood. From here on, we shall refer to this version as GravSphere1.5. Júlio et al. (2023) also introduced and made publicly available a version of GravSphere1.5 modified to test scalar field dark matter models in dwarf galaxies. [^5] Available at: https://github.com/marianajulio/alternative
@@
A version of GravSphere1.5 with an explicit treatment of rotation has also been presented in Júlio et al. (submitted). Lastly, self-consistent inclusion of millisecond pulsar accelerations as independent kinematic tracers to constrain mass-anisotropy models in star clusters and improved photometric modeling with an $\alpha\beta\gamma$ profile were also introduced in Bañares-Hernández et al. (2025), who used it to probe the presence of an IMBH and stellar remnant populations in Omega Centauri.
GravSphere has been a popular tool which has been used in the study of the dark matter distribution of nearby dwarf galaxies (e.g. Read et al. 2018, 2019; Gregory et al. 2019; Alvarez et al. 2020; Alvey et al. 2021; Collins et al. 2021; Zoutendijk et al. 2021; Júlio et al. 2023; De Leo et al. 2024; Zimmermann et al. 2025; Pickett et al. 2025), galaxy clusters (Li et al., 2023), and to model GCs (Bañares-Hernández et al., 2025).
In this section we introduce the improved methodology of GravSphere2 for mass-modeling stellar kinematics. We will briefly summarize the key elements of previous incarnations of GravSphere, focusing here on the novel aspects with respect to these earlier versions (Read & Steger, 2017; Read et al., 2018; Genina et al., 2020; Read et al., 2021; Collins et al., 2021). Crucially, these improvements include the removal of binning in favor of realistic self-consistent velocity distribution functions (Appendix A), the incorporation of the fourth-order Jeans equations to treat higher moments (with a more general treatment than previous approaches), and the inclusion of higher-order PMs. For details of mass-anisotropy models, see Appendix B.
GravSphere2 numerically solves the Jeans equations for a spherical, non-rotating [^6] As discussed in Bañares-Hernández et al. (2025), this assumption can be relaxed to some extent by solving for the second velocity moment directly, which we do. This is done by subtracting only global averages to the velocity distribution, leaving potential rotation signatures present. These are then reabsorbed when solving the Jeans equation, leading to a quadrature term that is implicitly added to the pure velocity dispersion (see e.g. the discussion in Fritz et al., 2016). system of tracers, with radial density profile, $\nu_{\star}(r)$, in dynamical pseudo-equilibrium (e.g. Binney & Tremaine, 2008):

### v2

@@ 1 Introduction
Many stellar systems are approximately spherical, including globular clusters (GCs) and nearby dwarf spheroidal galaxies (dSphs). Both GCs and dSphs contain seemingly invisible mass components that can be inferred from the anomalous motions of their stars and gas (see e.g. 1978PhDT.......195B; 1980ApJ...238..471R; 1987ARA&A..25..425T; 2002MNRAS.333..697L; van2010new; battaglia2013internal, for more recent examples in GCs and dSphs (respectively), see e.g. Evans:2021bsh; 2024MNRAS.529..331D; 2024ApJ...975..268S; 2025A&A...693A.104B and 2019MNRAS.484.1401R; 2020ApJ...904...45H; 2022NatAs...6..659B; 2023A&A...676A..63B; 2024MNRAS.535.1015D; 2025A&A...699A.347A; 2025A&A...700A..77P; 2025arXiv250820711V). For GCs, the invisible mass most likely constitutes stellar remnants or an ‘intermediate-mass black hole’ (IMBH), though some GCs may also contain dark matter. By contrast, dSphs appear to be dominated at all radii by dark matter. With increasingly available spectroscopic and astrometric data from recent surveys, GCs promise an unprecedented natural laboratory to study stellar remnants and hunt for IMBHs (2017IJMPD..2630021M; 2020A&ARv..28....4N; 2020ARA&A..58..257G; 2023arXiv231112118A; 2023MNRAS.522.5740V; 2024Natur.631..285H; 2025A&A...693A.104B), while dSphs promise the same for dark matter (2017arXiv170701348T; 2018MNRAS.481..860R; 2019MNRAS.484.1401R; 2022NatAs...6..659B; 2024MNRAS.535.1015D; 2024ApJ...970....1V; 2025arXiv250820711V; 2025PhRvL.134o1001Z).
The above proliferation in data has led to an associated proliferation in mass-modeling methodologies. All methods seek to solve the steady-state Collisionless Boltzmann Equation (CBE), also known as Vlasov Equation (e.g. Binney2008):
@@
By solving for the gravitational potential and subtracting the expected potential from all visible components we can infer, and study, the presence of invisible masses like black holes and dark matter (e.g. van2010new; 2009ApJ...704.1274W), which is usually done in the steady state approximation, assuming $\frac{\partial f}{\partial t}=0$.
@@
There are number of methods for solving the CBE. Jeans methods solve moments of the CBE (1922MNRAS..82..122J), having the advantages that they are fast [^1] This is particularly timely when assessing increasingly large samples with thousands (in some cases approaching millions, 2025ApJ...983...95H) of tracers. and minimize potential biases that could creep in due to assumptions about the form of the distribution function, $f$ (e.g. 2008MNRAS.390...71C; 2013MNRAS.436.2598W; 2013MNRAS.429.3079M; 2015arXiv150405533C; 2017MNRAS.471.4541R). The main disadvantages, however, are that the data typically need to be binned, while $f$ is not determined. This leads to some loss of information and makes it hard to incorporate stars of uncertain membership, binaries and/or rigorously model selection effects (e.g. 2021MNRAS.501..978R). Furthermore, since $f$ is not determined, a physical $f$ is not guaranteed. This can result in models that formally require a negative $f$ in some regions of phase space (1992ApJ...391..531D; 2013MNRAS.432.3361R) and that should, therefore, be discarded (e.g. 2006ApJ...642..752A).
@@
Distribution function methods assume some form for $f$ and then fit this to the data (e.g. 2015MNRAS.454..576G; 2019MNRAS.482.1525V; 2021MNRAS.501..978R; 2023MNRAS.522.5320D). $f$ can also be implicitly expressed as a superposition of orbits with ‘Schwarzschild methods’ (e.g. 1979ApJ...232..236S; 2008ApJ...682..841C; 2008MNRAS.385..647V; 2013MNRAS.433.3173B) or with ‘made-to-measure’ methods, in which $N$-body simulations are evolved to fit some data constraints (e.g. 1996MNRAS.282..223S; 2007MNRAS.376...71D; 2009MNRAS.395.1079D; 2013MNRAS.430.1928H), while 2025MNRAS.538.1442L propose a novel non-parametric distribution function method. Distribution function methods have the advantages that no data binning is required and $f$ is determined by the data. However, if the functional form for $f$ (or its implicit specification via other methods) does not encompass the true solution, then bias can creep in.
@@
Finally, numerous machine-learning methods have been proposed that learn from training data to either accelerate distribution function fitting or incorporate ‘cosmological priors’ into the mass-modeling pipeline (e.g. 2023MNRAS.519.4384E; 2025arXiv250303812N; 2025arXiv250500763L; 2025arXiv250307717S).
@@
Given the proliferation of mass-modeling methods, we may wonder why yet another method is needed. However, there are two key problems that remain to be addressed, even for spherical stellar systems. Firstly, while the amount and quality of data continue to improve, there will always be interesting systems with very few stellar tracers (e.g. 2021MNRAS.505.5686C; 2024ApJ...961...92S; 2025arXiv250504475P). Typically, simple mass estimators are used for low tracer numbers, but these have been shown to be potentially quite biased (2019MNRAS.484..245L). More sophisticated, unbiased, mass-modeling methods for low tracer numbers are only just being developed (e.g. 2025arXiv250303812N).
@@
Secondly, there is a well-known degeneracy between the cumulative mass (which we are typically most interested in) and the velocity anisotropy profile – a measure of how circular or radial the orbit distribution is (e.g. 1987MNRAS.224...13D; 1987ApJ...313..121M; 1990AJ.....99.1548M; Wilkinson:2001ut; 2003MNRAS.343..401L; 2009MNRAS.395...76D; 2017MNRAS.471.4541R; 2021MNRAS.501..978R). For Jeans modeling, this occurs because the Jeans equations do not constitute a closed system, meaning that for a given mass model and tracer profile, a unique solution is not guaranteed (e.g. Binney2008). This typically arises when only projected line-of-sight (LOS) velocities are available.
@@
Two key methods have been proposed to break the mass-anisotropy degeneracy. Firstly, one can incorporate higher moments of the velocity distribution as data constraints. At infinite order, these fully specify the phase-space distribution function and ensure its positivity (1990AJ.....99.1548M; 1992ApJ...391..531D; 2013MNRAS.432.3361R). After the usually-used second moment, the next non-trivial moment for a spherical, non-rotating, pressure-supported system is the fourth moment. This encodes information about deviations from Gaussianity, including the ‘tailedness’ of the distribution and the kurtosis. It can be incorporated directly into the Jeans equations, thereby introducing a new fourth-order anisotropy parameter (1990AJ.....99.1548M; 2003MNRAS.343..401L; 2005MNRAS.363..918L; 2013MNRAS.432.3361R; 2024arXiv240412671W). Alternatively, one can perform luminosity-weighted integrals over the fourth moment to derive ‘Virial Shape Parameters’ (VSPs) that marginalize out the higher-order anisotropy (1990AJ.....99.1548M; 2014MNRAS.441.1584R; 2017MNRAS.471.4541R). Secondly, one can include proper motions (PMs) to gain additional phase-space information that is independent of the line-of-sight velocities (e.g. van2010new; 2013MNRAS.436.2598W). Until recently, such PMs have only been available for GCs (e.g. 2003ApJ...595..187M; anderson2010new; 2022MNRAS.514..806V; 2023MNRAS.522.5320D; 2024Natur.631..285H; 2025A&A...693A.104B). However, data are becoming increasingly available also for nearby dwarf galaxies, with recent mass models incorporating PMs having been recently reported for the Draco and Sculptor dSphs (2024ApJ...970....1V; 2025arXiv250820711V).
@@
Although Jeans mass-modeling methods typically use data binning (e.g. 2017MNRAS.471.4541R), this is not actually a requirement. Several authors have explored assuming a local Gaussian distribution function constrained by the second moment to create a bin-free Jeans method (e.g. 2013MNRAS.436.2598W; 2013MNRAS.429.3079M). With the widespread availability of individually resolved stellar velocities, this approach is preferable and has all the advantages of having $f$ that distribution function methods enjoy. [^2] This is true, up to a given order of moments, provided that the Jeans method has those moments self-consistently implemented in the velocity distribution function.
@@
In this work, we present the GravSphere2 mass-modeling method. This builds on GravSphere(2017MNRAS.471.4541R; 2021MNRAS.505.5686C) with the specific goals of: (i) reducing bias; (ii) maximizing the information content in the data; (iii) avoiding the need for data binning; and (iv) being able to successfully model, without bias, very low numbers of tracers. To achieve these goals, we self-consistently solve the second and fourth-order Jeans equations, following the framework in 2013MNRAS.432.3361R and extending it to include a new observable: fourth-order PMs. We show that this breaks degeneracies in the higher-order Jeans equations by solving along their three projected components. Unlike previous approaches, which have made more restrictive assumptions (e.g. 1990AJ.....99.1548M; 2003MNRAS.343..401L; 2005MNRAS.363..918L; 2013MNRAS.432.3361R; 2024arXiv240412671W), we also include fully independent and radially varying anisotropies at both orders. We model individual stellar velocities, rather than binning the data, by introducing a local non-Gaussian velocity probability distribution function (PDF) constrained by the second and fourth moments. We also model the positions of stars without binning, where data are available. We test these improvements using mock data for dSphs and we make the Python package, GravSphere2, publicly available at: https://github.com/dadams42/GravSphere2.
@@
This paper is organized as follows. In Sec. 2, we introduce the modeling methodology behind GravSphere2, including our method for solving the fourth-order Jeans equations, which we apply following 2013MNRAS.432.3361R and extend to include PMs (Sec. 2.1), its implementation using non-Gaussian local velocity PDFs for individual tracers, following 2020MNRAS.499.5806S (Sec. 2.2), and the functional forms we assume for our mass model and the second and fourth-order stellar velocity anisotropy profiles (Sec. 2.3). In Secs. 3.1 and 3.2, we introduce the mock data which we use to test GravSphere2. In Sec. 3.3, we describe how we integrate the approach introduced in Sec. 2.1 to fit these data, specifying the form of the likelihood function and our assumed priors. Sec. 4 shows the results of our tests of GravSphere2 on the mock data. We compare these results with previous methods that have been applied to these same data in the literature, including previous versions of GravSphere. We also explore the limit of low tracer numbers (i.e. corresponding to the faintest or most distant systems) and we compare the performance of GravSphere2 with simple mass estimators that are widely used in the literature. In Sec. 5, we briefly discuss the significance of this work and we add some concluding remarks, highlighting some upcoming works that use the new GravSphere2 method.
@@ 2 Modeling stellar kinematics with GravSphere2
GravSphere is a numerical Jeans solver that was originally introduced in 2017MNRAS.471.4541R. It has since been further extended and extensively tested with mock data (2017MNRAS.471.4541R; 2018MNRAS.481..860R; 2019MNRAS.485.2010G; 2020MNRAS.498..144G; 2021MNRAS.501..978R; 2021MNRAS.505.5686C; 2022MNRAS.510.2186G; 2024MNRAS.535.1015D; 2025arXiv250303812N; 2025arXiv250418617T) including simulations deliberately chosen to explore expected bias in dynamical models due to breaking of key assumptions, such as spherical symmetry (triaxial mocks) and dynamical pseudo-equilibrium (tidally stripped mocks). A public version of this GravSphere1 code, called pyGravSphere, was introduced in 2020MNRAS.498..144G. [^3] Available at: https://github.com/AnnaGenina/pyGravSphere. Its most recent version was introduced and made publicly available in 2021MNRAS.505.5686C [^4] Available at: https://github.com/justinread/gravsphere. which improved the velocity binning with the binulator feature, improved the mass model and improved the way that the Virial Shape Parameters contributed to the likelihood. From here on, we shall refer to this version as GravSphere1.5. Julio:2023oyg also introduced and made publicly available a version of GravSphere1.5 modified to test scalar field dark matter models in dwarf galaxies. [^5] Available at: https://github.com/marianajulio/alternative
@@
A version of GravSphere1.5 with an explicit treatment of rotation has also been presented in 2025arXiv251006905J. Lastly, self-consistent inclusion of millisecond pulsar accelerations as independent kinematic tracers to constrain mass-anisotropy models in star clusters and improved photometric modeling with an $\alpha\beta\gamma$ profile were also introduced in 2025A&A...693A.104B, who used it to probe the presence of an IMBH and stellar remnant populations in Omega Centauri.
GravSphere has been a popular tool which has been used in the study of the dark matter distribution of nearby dwarf galaxies (e.g. 2018MNRAS.481..860R; 2019MNRAS.484.1401R; 2019MNRAS.485.2010G; 2020JCAP...09..004A; 2021MNRAS.501.1188A; 2021MNRAS.505.5686C; 2021arXiv211209374Z; Julio:2023oyg; 2024MNRAS.535.1015D; 2025PhRvL.134o1001Z; 2025arXiv250504475P; 2025arXiv251006905J), galaxy clusters (2023A&A...677A..24L), and to model GCs (2025A&A...693A.104B).
In this section we introduce the improved methodology of GravSphere2 for mass-modeling stellar kinematics. We will briefly summarize the key elements of previous incarnations of GravSphere, focusing here on the novel aspects with respect to these earlier versions (2017MNRAS.471.4541R; 2018MNRAS.481..860R; 2020MNRAS.498..144G; 2021MNRAS.501..978R; 2021MNRAS.505.5686C). Crucially, these improvements include the removal of binning in favor of realistic self-consistent velocity distribution functions (Sec. 2.2), the incorporation of the fourth-order Jeans equations to treat higher moments, with a more general treatment than previous approaches (Sec. 2.1), and the inclusion of higher-order PMs. Details of mass-anisotropy models are addressed in Sec. 2.3.
2.1 An extended Jeans analysis up to fourth order
GravSphere2 numerically solves the Jeans equations for a spherical, non-rotating system of tracers, with radial density profile, $\nu_{\star}(r)$, in dynamical pseudo-equilibrium (e.g. Binney2008):